Select all the expressions that are equivalent to 4 + w + 12 w .
A. 4 + 13 w
B. 13 w2 + 4
C. 2 (2 + 6w ) + w
D 16 + 2 w
E 2 ( 2 + 6w2)

Answer by Cromlix(4381)

Hi there,
4 + w + 12w
= 4 + 13w
A = 4 + 13w
C = 2(2 + 6w) + w
= 4 + 12w + w
= 4 + 13w
So A and C
Hope this helps :-)